marks each)

Item Quantity 

(2012)

Price 

(2012)

Price 

(2013)

Oranges 50 $0.90 $0.75

Bananas 100 $0.50 $0.95

Chicken 200 $2.00 $2.50

Beef 100 $5.00 $4.80

Bread 300 $1.75 $2.00

a) What is the cost of the CPI basket in 2012?

b) What is the cost of the CPI basket in 2013?

c) What is the CPI for 2012?

d) What is the CPI for 2013?


1
Expert's answer
2021-08-31T08:44:04-0400

Solution:

a.). The cost of CPI basket in 2012:

= (50 "\times" 0.90) + (100 "\times" 0.50) + (200 "\times" 2) + (100 "\times" 5) + (300 "\times" 1.75) = 45 + 50 + 400 + 500 + 525 = 1,520

 

b.). The cost of CPI basket in 2013:

= (50 "\times" 0.75) + (100 "\times" 0.95) + (200 "\times" 2.50) + (100 "\times" 4.80) + (300 "\times" 2.00) = 37.5 + 95 + 500 + 480 + 600 = 1,712.5


c.). What is the CPI for 2012:

CPI = "\frac{Cost \;in\;year}{Cost\;in \;base\;year} \times 100= \frac{1,520}{1,520} \times 100 = 100 \%"

 


d.). What is the CPI for 2013:

CPI = "\frac{Cost \;in\;year}{Cost\;in \;base\;year} \times 100= \frac{1,712.50}{1,520} \times 100 = 112.66 \%"